Grade 11Algebra

if one,omega and omega square are the cube root of unity the the solution of (x-3) the whole square+8=0 is

KEERTHANA
8 Years agoGrade 11

1 Answer

Adarsh
8 Years ago
Dear,
(x-3)2= -8
=>  x-3=2√2i          ( i = √-1  )
=> x= 3+ 2√2i
hope this will help ….
Thanks and Regards
Adarsh